# Conformal Convolution and Monte Carlo Meta-learners for Predictive Inference of Individual Treatment Effects

### v4

@@ 2 Background
Recently, Vovk et al. [17] proposed conformal predictive systems (CPS), an extension of full CP, that allow deriving predictive distributions and enjoys provable properties of validity, in the sense of probabilistic calibration, under IID assumption [17]. CPS produces conformal predictive distributions (CPD) as p-values arranged into a probability distribution function [17], created with the help of conformity measures. Vovk et al. [17] define a CPS as a function that is both a smooth conformal transducer and a randomized predictive system (RPS). The conformal transducer is another way to represent a conformal predictor, i.e., a confidence transducer that can be seen as a function that maps each input sequence into a sequence of p-values [14]. Formally, we define the full smoothed conformal transducer defined by conformity measure $A$ as,
$Q(z_{1},...,z_{n},(x,y),\phi):=\sum_{i=1}^{n+1}[R_{i}^{y}<R_{n+1}^{y}]\frac{1}{n+1}+\sum_{i=1}^{n+1}[R_{i}^{y}=R_{n+1}^{y}]\frac{\phi}{n+1}$
@@
where $(z_{1},...,z_{n})$ is the training sequence, $\phi\in[0,1]$, $x_{n+1}$ is a test object, and for each label $y$ the corresponding conformity score $R_{i}^{y}$ is defined as
$\begin{split}R_{i}^{y}&:=A(z_{1},...,z_{i-1},z_{i+1},...,z_{n},(x_{n+1},y),z_{i}),\qquad i=1,...,n\\ R_{n+1}^{y}&:=A(z_{1},...,z_{n},(x_{n+1},y)).\end{split}$
@@
The RPS is then a function $Q:\textbf{Z}^{n+1}\times[0,1]\rightarrow[0,1]$ that satisfies three requirements: (1) the function $Q(z_{1},...,z_{n},(x,y),\phi)$ needs to be monotonically increasing both in $y\in\mathbb{R}$ and $\phi\in[0,1]$; (2) $\lim_{y\rightarrow-\infty}Q(z_{1},...,z_{n},(x,y),0)=0$ and $\lim_{y\rightarrow\infty}Q(z_{1},...,z_{n},(x,y),1)=1$; and (3) it must be probabilistically calibrated, $\forall\alpha\in[0,1]:\mathbb{P}\{Q(z_{1},...,z_{n},z_{n+1},\tau)\leq\alpha\}=\alpha.$
A smoothed conformal transducer is an RPS if the IID assumption is satisfied and if it is defined by a monotonic conformity measure [17].
@@
CPS has been adapted to become more computationally efficient, i.e., split conformal predictive systems (SCPS) [27], by building upon ICP. Here, the p-values are created by a split conformity measure that is an element of a family of measurable functions $A_{m}:\{(X_{1},y_{1}),...,(X_{m},y_{m},(X,y)\}\rightarrow\mathbb{R}\cup\{-\infty,\infty\}$. This measure needs to be isotonic and balanced. Examples of such a measure are the residual error $y-\hat{f}(x)$ and normalized residual error $\frac{y-\hat{f}(x)}{\hat{u}(x)}$. The SCPS is a function that is both a split-smoothed conformal transducer and an RPS. The split conformal transducer determined by a split conformity measure $A$ is defined as
$Q(z_{1},...,z_{n},(x,y),\tau):=$ $\sum_{i=m+1}^{n}[R_{i}<R^{y}]\frac{1}{n-m+1}+\sum_{i=m+1}^{n}[R_{i}=R^{y}]\frac{\tau}{n-m+1}$
$+\frac{\tau}{n-m+1}$
where conformity scores $R_{i}$ and $R^{y}$ are defined by
$\begin{split}R_{i}&:=A(z_{1},...,z_{m},(x_{i},y_{i})),\qquad i=m+1,...,n,\\ R^{y}&:=A(z_{1},...,z_{m},(x,y)),\qquad y\in\mathbb{R}.\end{split}$
Vovk et al. [27] proved that the split smoothed conformal transducer is an RPS if the IID assumption is satisfied and the conformity measure is isotonic and balanced.
Conformal prediction and conformal predictive systems under covariate shift
@@
As mentioned, CP and CPS validity relies on the exchangeability and IID assumptions, respectively. However, in counterfactual inference, except in randomized control trials (RCTs), these assumptions are often violated due to covariate shift. Treatment assignment probabilities, depending on covariates, cause this shift, affecting the modeling of counterfactual outcomes.
@@
Fortunately, Tibshirani et al. [28] extended CP beyond the exchangeability assumption, proposing weighted CP (WCP), a weighted version of CP to deal with covariate shifts where the likelihood ratio between the training $P_{X}$ and test $\tilde{P}_{X}$ is known. WCP reweights each nonconformity score by a probability proportional to the likelihood ratio $\frac{d\tilde{P}_{X}(x_{i})}{dP_{X}(x_{i})}$. Lei and Candès [12] implemented it for counterfactual inference, next to extending it to CQR and giving an upper bound to the conservative validity guarantee of WCP. For CPS, Jonkers et al. [18] proposed weighted CPS (WCPS), similar to WCP, to allow for probabilistically calibrated predictive distributions under covariate shift.
@@ 3 Conformal Convolution and Conformal Monte Carlo Meta-Learner
The main concept behind both conformal convolution (CC) and conformal Monte Carlo (CMC) T-learner is to start from counterfactual inference. They aim to generate probabilistically calibrated predictive distributions to infer the ITE predictive distribution. The CCT- and CMC T-learner methods are illustrated in Figure 1, with descriptions provided in Algorithm 1 and 2.
Figure 1: Illustration of CCT- and CMC T-learner algorithms.
@@
2:  Input: training covariates $X_{tr}$, labels $y_{tr}$, treatment indicator $W_{tr}$, propensity score function $\pi(x)$, and test object $x_{test}$
3:  $X_{tr,0}$, $y_{tr,0}$:= $X_{tr}[W_{tr}=0]$, $y_{tr}[W_{tr}=0]$
4:  $X_{tr,1}$, $y_{tr,1}$:= $X_{tr}[W_{tr}=1]$, $y_{tr}[W_{tr}=1]$
5:  $\textbf{Q}_{Y^{0}}(x_{test},y,\phi):=$$\textbf{WCPS}(X_{tr,0}$, $y_{tr,0}$, $x_{test}$, $\frac{1}{1-\pi(x)}$,$y$,$\phi)$
6:  $\textbf{Q}_{Y^{1}}(x_{test},y,\phi):=$$\textbf{WCPS}(X_{tr,1}$, $y_{tr,1}$, $x_{test}$, $\frac{1}{\pi(x)}$, $y$, $\phi)$
7:  $\textbf{Q}_{ITE}(x_{test},y,\phi)=$$\textbf{Q}_{Y^{1}}*\textbf{Q}_{-Y^{0}}(x_{test},y,\phi)$see Eq. 6}
@@
Both methods deal with covariate shifts using WCPS. These shifts occur because the control and treatment models are trained and calibrated solely on observed outcomes, leading to different covariate distributions between treated and control groups due to confounding variables. To achieve calibrated predictive distributions for outcomes under control $Y^{0}$ and treatment $Y^{1}$, we weight each conformity scores proportional to $\frac{1}{1-\pi(x)}$ and $\frac{1}{\pi(x)}$ since
$w_{0}(x)$ $=\frac{dP_{X}(x)}{dP_{X|W=0}(x)}=\frac{dP_{X}(x)\mathbb{P}(W=0)}{dP_{X}(x)(1-\pi(x))}\propto\frac{1}{1-\pi(x)}$
$w_{1}(x)$ $=\frac{dP_{X}(x)}{dP_{X|W=1}(x)}=\frac{dP_{X}(x)\mathbb{P}(W=1)}{dP_{X}(x)\pi(x)}\propto\frac{1}{\pi(x)}$
@@
where $P_{X}(x)$ denotes the covariate distribution of the entire population, $P_{X|W=1}(x)$ and $P_{X|W=0}(x)$ represent the covariate distribution of the treated and control population, and $\mathbb{P}(W=1)$ and $\mathbb{P}(W=0)$ are the marginal probabilities of receiving treatment or not, respectively. These weights align with those used by Lei and Candès [12] for counterfactual inference with WCP and are similar to the inverse probability weighted (IPW) estimation of average causal effects [29].
@@
From the WCPS, we infer the potential outcomes predictive distributions, $Q_{Y^{0}}$ and $Q_{Y^{1}}$, which are probabilistically calibrated. Our interest lies in the ITE, $Y^{1}-Y^{0}$, so we perform a convolution operation on the treatment and the negated control predictive distributions, defined as:
[…]
This integral is intractable and requires numerical integration techniques, making the WCPS computationally expensive on top of the already expensive full WCPS. Therefore, we use a split WCPS approach in practice, defining the convolution as follows:
$\begin{split}Q_{ITE}(x_{test},\pi,y,\phi)&=\sum_{i=1}^{N_{cal,Y^{1}}}\sum_{j=1}^{N_{cal,Y^{0}}}[\{q_{i,Y^{1}}(x_{test})-q_{j,Y^{0}}(x_{test})\}<y]p_{i}^{w_{1}}(x_{test})p_{j}^{w_{0}}(x_{test})\\ &+[\{q_{i,Y^{1}}(x_{test})-q_{j,Y^{0}}(x_{test})\}=y]p_{i}^{w_{1}}(x_{test})p_{j}^{w_{0}}(x_{test})\phi\\ &+[1-\{1-p_{test}^{w_{1}}(x_{test})\}\{1-p_{test}^{w_{0}}(x_{test})\}]\phi\end{split}$ (6)
where
$p_{i}^{w}(x_{test})=\frac{w(x_{i})}{\sum_{j=1}^{n}w(x_{j})+w(x_{test})},\qquad p_{test}^{w}(x_{test})=\frac{w(x_{test})}{\sum_{j=1}^{n}w(x_{j})+w(x_{test})},$ (7)
$N_{cal,Y^{1}}$ and $N_{cal,Y^{0}}$ are the calibration set sizes for the split WCPS of outcomes under treatment and control, $q_{i,Y^{1}}(x_{test})$ and $q_{j,Y^{0}}(x_{test})$ are the $i$th and $j$th right-continuous points of $Q_{Y^{1}}(x_{test},y,\phi=0)$ and $Q_{Y^{0}}(x_{test},y,\phi=0)$, and $\pi$ represents the propensity score function.
While this split CCT-learner is way more computationally efficient, the convolution applied during inference, see Figure 1, is quite impractical in certain cases. Therefore, we propose the CMC T-learner, which replaces the convolution performed during inference with Monte Carlo sampling during training. The CMC framework samples from the fitted (potential) outcome distributions, creating Monte Carlo (MC) samples of the (potential) outcomes. Subsequently, these MC samples are utilized to artificially generate ITE samples. To generate these ITE samples, each example in the calibration set is sampled a specified number of times, producing multiple MC ITE samples for the same covariates. These ITE samples, treated as ground truth, are then employed alongside a meta-learner functioning as a CATE estimator to define a CPS, ultimately yielding conditional ITE predictive distributions.
We propose two methods for generating MC ITE samples: (1) The MC ITE method independently samples from both conditional potential outcome predictive distributions, subtracting the potential outcome under control $\hat{Y}^{0}$ from the one under treatment $\hat{Y}^{1}$; (2) The pseudo-MC (pMC) ITE method adapts this by conditioning on the observed treatment value, calculating the ITE sample as the difference between the observed outcome $Y$ and the conditional potential outcome predictive distribution under control $\hat{Y}^{0}$, and vice versa when treatment is not administered. Since the two approaches sample from different covariate distributions, they require a different weighting to be applied to WCPS. While the MC requires the same weighting scheme as in the CCT-learner, the pMC procedure requires a weighting of $\frac{\pi(x)}{1-\pi(x)}$ and $\frac{1-\pi(x)}{\pi(x)}$ for the counterfactual outcome of the treated and untreated subjects. Algorithm 3 provides a comprehensive overview of this sampling scheme.
We also introduce two variants of CMC meta-learners: CMC S- and X-learners, that follow the same process for generating conditional ITE predictive distributions, differing slightly in modeling nuisance predictive distributions and leveraging MC ITE samples. Details are provided in Appendix B.
Both CCT- and CMC meta-learners return conditional predictive distributions of the ITE. Prediction intervals of confidence level $1-\alpha$ are formed by taking the $\frac{\alpha}{2}$ and $1-\frac{\alpha}{2}$ quantiles from the predictive distribution as the lower and upper bounds, respectively. Our method produces symmetric, two-sided intervals with equally distributed probability density. In contrast, the approaches by Lei and Candès [12] and Alaa et al. [15] generate asymmetric intervals, offering an efficiency advantage due to smaller interval lengths [22]. However, these asymmetric intervals can be misleading in practical applications as they might intuitively imply equal probability density on both sides of the interval.
3.1 The Epsilon Problem
In addition to introducing the CCT- and CMC meta-learners, we propose to add a fourth assumption to the three assumptions (see Section 2) on the underlying data-generating processes to perform CATE estimation to define its relation with the validity of predictive inference on ITEs. This fourth assumption covers the relation between noise distributions under treatment $\varepsilon(1)$ and control $\varepsilon(0)$. If we want to use the CP framework for performing valid predictive inference, we need to know their joint distribution. As we never jointly observe $\varepsilon(0)$ and $\varepsilon(1)$, and can thus not learn their joint distribution, we need to assume how these two noise distributions are related. The literature related to CATE estimation often assumes, when creating data-generating processes, that these noise terms are zero-mean random variables and independent of X and W  [4, 5, 6, 7]. Some works [5, 6, 7] assume that the treatment effect is deterministic and thus, in consequence, that the two noise terms are identical for the same example, whereas Künzel et al. [4] assumes that the noise distributions are fully independent in their synthetic simulation experiments. The limited literature concerning predictive inference of ITE  [10, 11, 12, 15] mostly assume that the noise distributions are fully independent and set up their experiments in the same way. In the following paragraphs, we will therefore discuss three possible characteristics of the noise terms’ joint distribution and their impact on the validity of predictive intervals and distributions of ITE. While the true nature of these joint noise distributions is, in part, a philosophical question, it requires further attention from the research community and domain experts.
$\mathbf{\boldsymbol{\varepsilon(0)=\varepsilon(1)}}$: In the scenario where the two noise terms are identical for each sample, a common assumption in most CATE simulation experiments, it is essential to acknowledge that the noise distributions become nullified when inferring ITE. This way, the ITE becomes entirely deterministic. Here both the CCT- and CMC meta-learners over account for uncertainty since the noise distribution (i.e., the aleatoric part) is shared in both (potential) outcomes. To address this, we suggest modeling the epistemic uncertainty distribution of two potential outcome models as future work. Assuming these epistemic uncertainty distributions are independent of each other, the convolution and the MC sampling become more valid. Nevertheless, it is important to note that here the CCT- and CMC meta-learners will continue to provide conservatively valid prediction intervals.
[…]
In this paper, simulation experiments, presented in Figure 3, further support these statements. Note that these assumptions also affect other predictive inference approaches for the ITE and are, therefore, not limited to CCT- and CMC alone.
3.2 Universal Consistent CCT- and CMC T-Learner
The introduction of the fourth assumption allows us to prove the existence of a universal consistent CCT- and CMC T-learner, i.e., the conformal predictive distribution approaches the true data-generating distribution as $n\rightarrow\infty$ under any date-generating distribution, which we formalize in Theorem 1.

### v5

@@ 3 Conformal predictive systems
Vovk et al. (2019, 2020) propose conformal predictive systems (CPS), an extension of CP (Appendix A.1) that allows deriving predictive distributions and enjoys provable properties of validity, in the sense of probabilistic calibration, under exchangeability (Vovk et al., 2019). CPS produces conformal predictive distributions as p-values arranged into a probability distribution function Vovk et al. (2019), created with the help of conformity scores. Vovk et al. (2019) define a CPS as a function that is both a smooth conformal transducer and a randomized predictive system (RPS). The conformal transducer is another way to represent a conformal predictor, i.e., a confidence transducer that can be seen as a function that maps each input sequence into a sequence of p-values (Vovk et al., 2022). Formally, we define the full smoothed conformal transducer, defined by conformity score $s$, as
$Q(X_{n+1},y,\phi;Z_{1:n}):=\sum_{i=1}^{n+1}[S_{i}^{y}<S_{n+1}^{y}]\frac{1}{n+1}+\sum_{i=1}^{n+1}[S_{i}^{y}=S_{n+1}^{y}]\frac{\phi}{n+1}$
@@
where $Z_{1:n}$ is the training sequence, $\phi\sim Uniform(0,1)$, $X_{n+1}$ is a test object, and for each label $y$ the corresponding conformity score $S_{i}^{y}$ is defined as
$\begin{split}S_{i}^{y}&:=s((X_{i},Y_{i});Z_{1:n\setminus i}\cup(X_{n+1},y)),\qquad i=1,...,n\\ S_{n+1}^{y}&:=s((X_{n+1},y);Z_{1:n}).\end{split}$
@@
An RPS is defined as a function $Q:\mathcal{Z}^{n+1}\times[0,1]\rightarrow[0,1]$ that satisfies three requirements: (1) the function $Q(X_{n+1},y,\phi;Z_{1:n})$ needs to be monotonically increasing both in $y\in\mathbb{R}$ and $\phi\in[0,1]$; (2) $\lim_{y\rightarrow-\infty}Q(X_{n+1},y,0;Z_{1:n})=0$ and $\lim_{y\rightarrow\infty}Q(X_{n+1},y,1;Z_{1:n})=1$; and (3) it must be probabilistically calibrated, $\forall\alpha\in[0,1]:\mathbb{P}\{Q(X_{n+1},Y_{n+1},\phi;Z_{1:n})\leq\alpha\}=\alpha.$
@@
A smoothed conformal transducer is an RPS if and only if the exchangeability assumption is satisfied and if it is defined by a monotonic conformity score (Vovk et al., 2019). CPS has been adapted to become more computationally efficient, i.e., split conformal predictive systems (SCPS) (Vovk et al., 2020), by building upon inductive CP (ICP). Here, the p-values are created by a split conformity score function that is an element of a family of measurable functions $s:\mathcal{Z}\times\mathcal{Z}^{m}\rightarrow\mathbb{R}\cup\{-\infty,\infty\}$, where $m$ is the length of the training sequence. This score needs to be isotonic and balanced. Examples of such a measure are the residual error $y-\hat{f}(x)$ and normalized residual error $\frac{y-\hat{f}(x)}{\hat{u}(x)}$, where $\hat{f}(\cdot)$ represent the fitted regression model and $\hat{u}(\cdot)$ a 1D-uncertainty estimate. The SCPS is a function that is both a split-smoothed conformal transducer and an RPS and follows a similar form as in Eq. 3, however, only leveraging conformity scores of the calibration set. Vovk et al. (2020) proved that the split smoothed conformal transducer is an RPS if the exchangeability assumption is satisfied and the conformity measure is isotonic and balanced. In this work, we also introduce a new conformity score for CPS, namely, the probability integral transform (PIT)-values $s=\hat{F}_{X_{i}}(Y_{i})$ of the distribution prediction $\hat{F}_{X_{i}}$. For this score, you need access to the CDF estimate and thus require a distributional regression approach as a base-learner. Note that Chernozhukov et al. (2021) introduced a similar score in their distributional CP approach.
3.1 Conformal predictive systems under covariate shift
@@
CP and CPS validity rely on the exchangeability assumption. However, in counterfactual inference, except in randomized control trials (RCTs), these assumptions are often violated due to covariate shifts. Treatment assignment probabilities, depending on covariates, cause this shift, affecting the modeling of counterfactual outcomes.
@@
Fortunately, Tibshirani et al. (2019) extended CP beyond the exchangeability assumption, proposing weighted CP (WCP), a weighted version of CP to deal with covariate shifts where the likelihood ratio between the training $\mathcal{P}_{X}$ and test $\tilde{\mathcal{P}}_{X}$ is known. WCP reweights each nonconformity score by a probability weight $p^{w}_{i}$ proportional to the likelihood ratio $w_{i}=\frac{d\tilde{\mathcal{P}}_{X}(X_{i})}{d\mathcal{P}_{X}(X_{i})}$, where the weights are defined by $p^{w}_{i}=\frac{w(X_{i})}{\sum^{n+1}_{j=1}w(X_{j})}$. Lei and Candès (2021) implemented WCP for counterfactual inference, next to extending it to conformalized quantile regression (CQR) and giving an upper bound to the conservative validity guarantee of WCP. For CPS, Jonkers et al. (2024) proposed weighted CPS (WCPS), similar to WCP, to allow for probabilistically calibrated predictive distributions under covariate shift. However, while Jonkers et al. (2024) proposes the framework for WCPS by introducing a weighted (split) conformal transducer, they omit the proof of the probabilistic calibration of the weighted conformal transducer (Theorem 1) and only validate it empirically. We introduce this proof in this work (Appendix E.1).
Theorem 1.
Assume that $Z_{i}=(X_{i},Y_{i})\in\mathcal{X}\times\mathbb{R}$, $i=1,...,n$ are produced independently from $\mathcal{P}=\mathcal{P}_{X}\times\mathcal{P}_{Y|X}$; $Z_{n+1}=(X_{n+1},Y_{n+1})\in\mathcal{X}\times\mathbb{R}$, is independently drawn from $\tilde{\mathcal{P}}=\tilde{\mathcal{P}}_{X}\times\mathcal{P}_{Y|X}$; $\tilde{\mathcal{P}}_{X}$ is absolutely continuous with respect to $\mathcal{P}_{X}$; likelihood ratio $w(x)=\frac{d\tilde{\mathcal{P}}_{X}}{d\mathcal{P}_{X}}$ representing the covariate shift, random number $\phi\sim Uniform(0,1)$; $Z_{1:n}$, $Z_{n+1}$, and $\phi$ to be independent. Then, the distribution of the weighted conformal transducer, defined by
$Q(X_{n+1},y,\phi,w(x);Z_{1:n}):=\sum_{i=1}^{n+1}[S_{i}^{y}<S_{n+1}^{y}]p_{i}^{w}+\sum_{i=1}^{n+1}[S_{i}^{y}=S_{n+1}^{y}]p_{i}^{w}\phi$ (1)
is uniform:
$\forall\alpha\in[0,1]:\mathbb{P}_{Z_{1:n}\sim\mathcal{P},(X_{n+1},Y_{n+1})\sim\tilde{\mathcal{P}}}\{Q(X_{n+1},Y_{n+1},\phi,w(x);Z_{1:n})\leq\alpha\}=\alpha$ (2)
This result allows us to generate potential outcome predictive distributions, which are guaranteed to be probabilistically calibrated in a finite sample setting under a known covariate shift, for both binary and continuous interventions. Theorem 1 also applies to split weighted conformal transducer and thus to split WCPS, as this is just a special case of WCPS. In the remaining part of this work, we will work with split WCPS, as it is more computationally efficient and easier to validate whether the score function is isotonic, which is necessary to generate a valid predictive distribution. Note that the approach of Jonkers et al. (2024) is model-agnostic; by building our framework upon WCPS, it can be used with any point and distributional regression model. We even show in our experiments (Section 5.2) that we can wrap WCPS and the in this paper presented distributional meta-learners (Section 4) around the flexible causal distributional models (FCCN (Zhou et al., 2022) and NOFLITE (Vanderschueren et al., 2023)). In this case, you could view WCPS as a hedge against miscalibration, at the cost of fewer samples to train the model, as some are needed for the calibration set. Additionally, Theorem 1 also allows us to obtain probabilistically calibrated distributions of individual treatment effects under treatment (ITT) and control (ITC), which can be used as an analysis tool.
@@ 4 Conformal convolution and conformal Monte Carlo meta-learner
We introduce the conformal convolution T- (CCT) and conformal Monte Carlo (CMC) meta-learners that both start from counterfactual inference. They use WCPS to generate probabilistically calibrated predictive distributions of the potential outcomes (Section 3) to infer the ITE predictive distribution. The CCT- and CMC meta-learner methods are illustrated in Figure 1, with descriptions in Algorithm 1 and 2.
Figure 1: Illustration of CCT- and CMC T-learner algorithms, where $\frac{1}{1-\pi}$ and $\frac{1}{\pi}$ are the proportional likelihood ratio function used to unconfound the outcome under control $Y(0)$ and treatment $Y(1)$, respectively.
@@
2:  Input: observations $\{Z_{i}=(Y_{i},X_{i},W_{i})\}_{i=1}^{n}$, propensity score function $\pi(x)$, and test object $X_{n+1}$.
3:  $Z^{W=0}_{1:n_{0}}:=\{(Y_{i},X_{i})|W_{i}=0\}_{i=1}^{n}$ where $n_{0}$ the number observation under control.
4:  $Z^{W=1}_{1:n_{1}}:=\{(Y_{i},X_{i})|W_{i}=1\}_{i=1}^{n}$ where $n_{1}$ the number observation under treatment.
5:  Construct $\textbf{Q}_{Y(0)}(X_{n+1},y,\phi_{0},w_{0}=\frac{1}{1-\pi};Z^{W=0}_{1:n_{0}})$ and $\textbf{Q}_{Y(1)}(X_{n+1},y,\phi_{1},w_{1}=\frac{1}{\pi};Z^{W=1}_{1:n_{1}})$.
6:  $\textbf{Q}_{ITE}(X_{n+1},y,\phi;\textbf{Q}_{Y(0)},\textbf{Q}_{Y(1)}))=\textbf{Q}_{Y(1)}*\textbf{Q}_{-Y(0)}(X_{n+1},y,\phi)$Eq. 3}
Algorithm 2 Fit and inference of CMC meta-learner
1:  Assume: $\phi\sim Uniform(0,1)$, $y\in\mathbb{R}$
@@
2:  Input: observations $\{Z_{i}=(Y_{i},X_{i},W_{i})\}_{i=1}^{n}$, propensity score function $\pi(x)$, test object $X_{n+1}$, and MC samples $n_{MC}$.
3:  $Z_{train,1:n_{train}},Z_{cal,1:n_{cal}}=\textbf{RandomSplit}(Z_{1:n})$
4:  $Z^{W=0}_{train,1:n_{0}}:=\{(Y_{train,i},X_{train,i})|W_{train,i}=0\}_{i=1}^{n_{train}}$ where $n_{0}$ the number observation in the training split under control.
5:  $Z^{W=1}_{train,1:n_{1}}:=\{(Y_{train,i},X_{train,i})|W_{train,i}=1\}_{i=1}^{n_{train}}$ where $n_{1}$ the number observation in the training split under treatment.
6:  Construct $\textbf{Q}_{Y(0)}(x,y,\phi_{0},w_{0}=\frac{1}{1-\pi};Z^{W=0}_{train,1:n_{0}})$ and $\textbf{Q}_{Y(1)}(x,y,\phi_{1},w_{1}=\frac{1}{\pi};Z^{W=1}_{train,1:n_{1}})$.
7:  $ITE_{MC,1:n_{ITE}},X_{MC,1:n_{ITE}}=\textbf{MonteCarloITE}(Z_{cal,1:n_{cal}},\textbf{Q}_{Y(0)},\textbf{Q}_{Y(1)},n_{MC})$ where $n_{ITE}=n_{cal}n_{MC}.$Algorithm 3}
8:  Construct $\textbf{Q}_{ITE}(X_{n+1},y,\phi;\{ITE_{MC,i},X_{MC,i}\}_{i=1}^{n_{ITE}})$.
@@
Both approaches deal with covariate shifts using WCPS. These shifts occur because the control and treatment models are trained and calibrated solely on observed outcomes, leading to different covariate distributions between treated and control groups due to confounding variables. To achieve calibrated predictive distributions for outcomes under control $Y(0)$ and treatment $Y(1)$, respectively, we weight each conformity score proportional to $\frac{1}{1-\pi(x)}$ and $\frac{1}{\pi(x)}$ since
$w_{0}(x)$ $=\frac{d\mathcal{P}_{X}(x)}{d\mathcal{P}_{X|W=0}(x)}=\frac{d\mathcal{P}_{X}(x)\mathbb{P}(W=0)}{d\mathcal{P}_{X}(x)(1-\pi(x))}\propto\frac{1}{1-\pi(x)}$
$w_{1}(x)$ $=\frac{d\mathcal{P}_{X}(x)}{d\mathcal{P}_{X|W=1}(x)}=\frac{d\mathcal{P}_{X}(x)\mathbb{P}(W=1)}{d\mathcal{P}_{X}(x)\pi(x)}\propto\frac{1}{\pi(x)}$
@@
where $\mathcal{P}_{X}(x)$ denotes the covariate distribution of the entire population, $\mathcal{P}_{X|W=1}(x)$ and $\mathcal{P}_{X|W=0}(x)$ represent the covariate distribution of the treated and control population, respectively, and $\mathbb{P}(W=1)$ and $\mathbb{P}(W=0)$ are the marginal probabilities of receiving treatment and control. The weights match those used by Lei and Candès (2021) for counterfactual inference with WCP and are similar to the inverse probability weighted (IPW) estimation of average causal effects (Imbens and Rubin, 2015).
@@
From WCPS, we infer the potential outcomes predictive distributions, $Q_{-Y(0)}$ and $Q_{Y(1)}$, which are probabilistically calibrated. Since our primary interest often lies in ITE, defined as $ITE=Y(1)-Y(0)$, we introduce a convolution-like operation on the treatment and the negated control predictive distributions, as defined in Eq. 3. Theorem 2 states the finite-sample calibration guarantee of the CCT-learner, and the proof is given in Appendix E.2.
@@
Assume that $Z_{i}=(X_{i},W_{i},Y_{i})\in\mathcal{X}\times\{0,1\}\times\mathbb{R}$, $i=1,...,n$ and $X_{n+1}\in\mathcal{X}$, $Y_{n+1}(0),Y_{n+1}(1)\in\mathbb{R}$ are produced independently from $\mathcal{P}$, and $\phi\sim Uniform(0,1)$.
Then, the "convolution" of the distributions $Q_{-Y(0)}$ and $Q_{Y(1)}$, defined by
$\begin{split}Q_{Y(1)-Y(0)}(X_{n+1},y,\phi,w_{0},w_{1};Z_{1:n})&=\sum_{i=1}^{n_{0}}\sum_{j=1}^{n_{1}}\left[\left\{q_{Y(1),j}(X_{n+1})-q_{Y(0),i}(X_{n+1})\right\}<y\right]p_{j}^{w_{1}}p_{i}^{w_{0}}\\ &+\left[\left\{q_{Y(1),j}(X_{n+1})-q_{Y(0),i}(X_{n+1})\right\}=y\right]p_{j}^{w_{1}}p_{i}^{w_{0}}\phi\\ &+\left[1-\left\{1-p_{n+1}^{w_{1}}\right\}\left\{1-p_{n+1}^{w_{0}}\right\}\right]\phi,\end{split}$ (3)
where $n_{1}$ and $n_{0}$ are the calibration set sizes for the WCPS of outcomes under treatment and control, $q_{Y(1),j}(X_{n+1})$ and $q_{Y(0),i}(X_{n+1})$ are the $j$th and $i$th right-continuous points of $Q_{Y(1)}(X_{n+1},y,\phi_{1}=0,w_{1}=\frac{1}{\pi};Z^{W=1}_{1:n_{1}})$ and $Q_{Y(0)}(X_{n+1},y,\phi_{0}=0,w_{0}=\frac{1}{1-\pi};Z^{W=0}_{1:n_{0}})$, is uniform:
$\begin{split}\forall\alpha\in[0,1]:&\mathbb{P}_{Z_{1:n},X_{n+1},Y_{n+1}(0),Y_{n+1}(1)\sim\mathcal{P}}\{\\ &Q_{Y(1)-Y(0)}(X_{n+1},Y_{n+1}(1)-Y_{n+1}(0),\phi,w_{0},w_{1};Z_{1:n})\leq\alpha\}=\alpha\end{split}$ (4)
when we assume weighted exchangeability according to,
$Y_{n+1}(1)-Y_{n+1}(0)|E_{q_{Y(1)}-q_{Y(0)}}\sim\sum_{i=1}^{n_{0}}\sum_{j=1}^{n_{1}}p^{w_{0}}_{i}p^{w_{j}}_{i}\delta_{q_{Y(1),j}-q_{Y(0),i}}+[1-\{1-p_{n+1}^{w_{1}}\}\{1-p_{n+1}^{w_{0}}\}]\delta_{\infty}$ (5)
where $E_{ITE}$ denotes the event that $\{\Psi_{1},\ldots,\Psi_{n0*n1+1}\}=\{\psi_{1},\ldots,\psi_{n0*n1+1}\}$, $\Psi_{i}=q_{Y(1),i\mathbin{//}n1}-q_{Y(0),i\mathbin{\%}n0}$ for $i\in[1,n]$, and $\Psi_{n0*n1+1}=Y_{n+1}(1)-Y_{n+1}(0)=ITE_{n+1}$.
While the CCT-learner comes with a finite sample guarantee, the convolution applied during inference (Figure 1) is impractical in some instances. Therefore, we propose the CMC meta-learner, which performs Monte Carlo (MC) sampling during training (and, optionally, fits an additional model). Specifically, we draw samples from each estimated potential outcome distribution to create MC replicates, then use these replicates to produce ITE samples. For each calibration example, we repeat this procedure multiple times, treating the resulting MC ITEs as pseudo ITEs, and train a CATE meta-learner on these samples to form a calibrated predictive distribution of ITEs. Depending on how we incorporate treatment assignment, we recover the S-, T-, and X-learner variants: the S-learner augments the covariates with treatment, the T-learner fits separate outcome models, and the X-learner additionally refits a model on the MC ITE samples. Note that this reduced computation at inference time comes at a cost as the CMC-learner requires two calibration sets.
Epsilon problem
The "Epsilon Problem" highlights a fundamental challenge in the predictive inference of ITE: the joint distribution of counterfactual noise terms, $\epsilon(0)$ and $\epsilon(1)$, is unidentifiable as they are never observed simultaneously. Existing approaches implicitly assume independent noise across treatments, or an identical relation, assumptions that are untestable yet critical for valid predictive distributions. We formalize that the ITE distribution is only identifiable if potential outcome marginals are independent or linked by a specified copula (Proposition 1). Violations of independence—such as correlated noise due to shared confounders—compromise distributional validity, though conservatively valid prediction intervals remain achievable. Simulations (Figure 11) empirically demonstrate coverage degradation under noise dependency. While partial identification via bounds offers an alternative, we pragmatically adopt the independent copula assumption (making a fourth assumption), aligning with formal causal models, to enable actionable uncertainty quantification despite its philosophical limitations. In Appendix D, we further elaborate on this problem.
Universal consistent CCT- and CMC T-learner
The introduction of the fourth assumption allows us to prove the existence of a universal consistent CCT- and CMC T-learner, i.e., the conformal predictive distribution approaches the true data-generating distribution as $n\rightarrow\infty$ under any data generating distribution, which we formalize in Theorem 3.
Theorem 3.